OpenStudy (anonymous):

find h'(5) when h(x)=10u(x)v(x) when u(5)=7, u'(5)=-3, v(5)=2, and v'(5)=6

OpenStudy (anonymous):

Oh these problems. First find h'(x) by using the product rule: h'(x)=10(u'(x)v(x)+u(x)v'(x)) Then evaluate for x=5

OpenStudy (anonymous):

do i take the derivative on each the u and v or take the regular u and v?

OpenStudy (anonymous):

It depends. The product rule tells you which one. If you look at what i wrote, some have a ' which means take the derivative. So, h'(5)=10(u'(5)v(5)+u(5)v'(5)) They give you all the values to the right so all you need to do is substitute them in.

OpenStudy (anonymous):

That makes sense. Thank you once again